OpenStudy (anonymous):

Using the Pythagorean theorem, find the measure of the hypotenuse when a = 7.2m and b = 5.5m. A. 6m B. 9m C. 12m D. 13m

OpenStudy (anonymous):

a^2+b^2=c^2 so (7.2)^2+(5.5)^2=c^2 82.09=c^2 find the square root 9.060

OpenStudy (anonymous):

a^2 + b^2 = c^2 c = sqrt( a^2 + b^2) c = sqrt( 7.2^2 + 5.5^2) solve

OpenStudy (anonymous):

do \[a^2+b^2=c^2\]

OpenStudy (anonymous):

a=7.2 b=5.5 now plug in!